OpenStudy (anonymous):

Find the length of a rectangle if the width is 8cm and the perimeter 40cm. A 24cm B 32cm C 5cm D 12cm

OpenStudy (a1234):

16 + 2l = 40 2l = 24 l = 12

OpenStudy (anonymous):

well perimeter of a square/rectangle P=a+a+a+a P = l + l+ w + w you already have width